OpenStudy (anonymous):

What is the y-intercept of the function y = 16x? 16 1 y x

OpenStudy (agent47):

y intercept means x=0 when x=0 y=16*0=0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

it woul be ZERO like what other people said..

OpenStudy (anonymous):

Im confused cause my answer key have no zeros would it be just x?

OpenStudy (anonymous):

 this is graph of your function

OpenStudy (anonymous):

@Agent47 just showed u..

OpenStudy (anonymous):

@myko thank u mybe he/she understand

OpenStudy (anonymous):

non of the answers in answer key is right tho

jimthompson5910 (jim_thompson5910):

The equation is probably y = 16^x, so the y-intercept is (0,1) because you get y = 1 when you plug in x = 0

OpenStudy (anonymous):

maybe like this: x-intercept=y-intercept

OpenStudy (anonymous):

you get 0 when you plug in x=0

OpenStudy (anonymous):

is you equation \[y = 16^x\]????

OpenStudy (anonymous):

@Suckatmath94

OpenStudy (anonymous):

so its 16 cause thats what i got?

jimthompson5910 (jim_thompson5910):

What is 16^0 equal to?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yeah i got 1